How does one fined the p-value from ANOVA analysis using a table? Use the example data below to derive a p-value.

treatment1 treatment2 treatment3
595 639 417
580 615 449
508 511 517
583 573 438
633 648 415
517 677 555
sum 3416 3663 2791
average 569.333333 610.5 465.166667
grand mean 548.333333
Mstr 33,667.20
MSE 3071.6
F 10.96
Fstat 3.68
Answer #1

k=3

n=3*6=18

Degree of freedom for between group = k - 1 = 3-1 =2

Degree of freedom for within group = n - k = 18 - 3 = 15

P value = 0.0012................by using Excel command =FDIST(10.96,2,15)

We can directly use here one way anova by Excel.

Step 1 ) Enter data in Excel.

Step 2) Data >>Data analysis >> One way anova >>Select data >>click on label in first row >>ok

Anova: Single Factor 633 4491 lalalalala SUMMARY Groups Count 6 6 Average Vartance 16 A 0.667 36 6105 93.5 2 4 .1057 333837 t